If α=3i−j+2k, , β=2i+j−k and γÌ…=i−2j+2k, find α×β×γ.
a.24i+7j-5k
b.-24i+7j-5k
c.24i+7j+5k
d.24i-7j-5k
1
Expert's answer
2020-06-09T16:15:03-0400

"\alpha X \beta =\\\begin{vmatrix}
 i & j & k \\
 3 & -1 & 2 \\
 2 & 1 & -1
\end{vmatrix}"


= i(1-2)-j(-3-4)+k(3+2)


= -i +7j +5k


"\alpha X \beta X \gamma =\\\begin{vmatrix}
 i & j & k \\
 -1 & 7 & 5 \\
 1 & -2 & 2
\end{vmatrix}"


= i(14+10) -j(-2-5) +k (2-7)


= 24i + 7j -5k

Option A is correct